# MSDformer: Multi-scale Discrete Transformer For Time Series Generation

### v1

@@ V-A Theoretical Analysis of Discrete Token Modeling
For a stationary ergodic source $X\sim P_{X}$ and a bounded distortion measure $d:\mathcal{X}\times\tilde{\mathcal{X}}\to\mathbb{R}^{+}$, the minimum achievable rate $R(D)$ at distortion level $D$ is given by:
$R(D)=\inf_{\begin{subarray}{c}p(\tilde{X}|X):\mathbb{E}[d(X,\tilde{X})]\leq D\end{subarray}}I(X;\tilde{X}),$ (23)
where $I(X;\tilde{X})$ is the mutual information between the source $X$ and the reconstruction $\tilde{X}$.
For the time series generation task, we define the data format as $x_{1:\tau}=(x_{1},\ldots,x_{\tau})\in\mathbb{R}^{\tau\times d}$, and $\tilde{x}_{1:\tau}$ represents the reconstructed data. The distortion measure is defined as the Mean Squared Error:
@@
In the vector quantization method, the time series data $x_{1:\tau}$ is compressed into a discrete token sequence $y_{1:L}\in\{0,1,\ldots,V-1\}^{L}$. The rate-distortion control for this method is described in Theorem 2.
@@
$\mathbb{E}[d(x_{1:\tau},\tilde{x}_{1:\tau})]\leq D.$ (27)
@@
V-B Theoretical Analysis of Multi-scale Modeling
In Section V-A, we conclude that increasing the codebook size directly raises the rate, which in turn allows for lower distortion. In MSDformer, the multi-scale modeling, which involves learning multi-scale discrete token representations of time series, indirectly increases the codebook size. To compare the merits of indirectly increasing the codebook size by adding scales versus directly increasing the codebook size on a single scale, this part will conduct a comparative analysis from a theoretical perspective.
@@ V-B Theoretical Analysis of Multi-scale Modeling
[…]
@@
[…]
In our experimental setup, the ratios $\frac{V^{\prime}}{V}$ and $\frac{L^{\prime}}{L}$ are constrained to the interval $\left[\frac{1}{4},4\right]$, and they share the same directional relationship (both are either $\geq 1$ or $\leq 1$). Furthermore, $V$ and $V^{\prime}$ are set within the range $[2^{7},2^{10}]$. Under these conditions, it is observed that $L\log_{2}\left(1+\frac{V^{\prime}}{V}\right)<L^{\prime}\log_{2}V^{\prime}$, which leads to the conclusion that $R_{m}>R_{s}$ always holds true. Therefore, in general, employing multi-scale modeling is more effective in enhancing the rate compared to directly increasing the codebook size on a single scale, thereby allowing for lower distortion. This theoretically validates the rationality of the MSDformer design.
VI Experiments
@@
VI-A Datasets and Metrics
We evaluate the proposed method on six benchmark datasets, including four real-world datasets (Stocks, ETT, Energy, and fMRI) and two simulated datasets (Sines and MuJoCo).
To quantitatively evaluate the synthetic data, we employ three metrics: 1) The Discriminative Score assesses distributional similarity by measuring how well a classifier can distinguish between real and synthetic samples; 2) The Predictive Score evaluates the practical utility of synthetic data by evaluating the performance of models trained on synthetic samples when tested against real data; 3) The Context-FID Score quantifies sample fidelity through representation learning, specifically comparing the distribution of local temporal features between real and synthetic time series.
@@ VI-A Datasets and Metrics
VI-B Short-term Time Series Generation
TABLE I: Results of all methods on all datasets
@@ VI-B Short-term Time Series Generation
TABLE II: Results of long-term time series generation
@@
Table I summarizes the performance of each algorithm across all datasets. The results in bold indicate the best performance, while the results with an underline indicate the second-best performance. These results allow us to draw several key observations.
@@
Fig. 3 and Fig. 4 show the comparative visualization results of MSDformer, SDformer, and Diffusion-TS across all datasets. These visualizations depict the distribution of synthesized and real time series data in a two-dimensional space and their probability density distributions. The results indicate that the data generated by MSDformer and SDformer closely resemble real data, highlighting the superior performance of DTM-based methods in time series generation tasks. Furthermore, through a detailed comparison, it can still be observed that the performance of MSDformer is better than that of SDformer. For example, in the Stock dataset, when the x-axis value is around 0.7, the kernel density estimation of the time series data generated by SDformer has a slight deviation from the kernel density estimation of the real data, while MSDformer is significantly closer to the real data at this position. This further illustrates the advantages of multi-scale modeling.
VI-C Long-term Time Series Generation
@@
MSDformer
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
((a))
((b))
((c))
((d))
((e))
((f))
SDformer
Diffusion-TS
MSDformer
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
((a))
((b))
((c))
((d))
((e))
((f))
SDformer
Diffusion-TS
VI-D Efficiency Analysis
@@
SDformer 0.15 0.18 0.40 0.15 0.15 0.44
MSDformer 0.95 0.96 0.30 0.97 0.96 1.40
To validate the practical efficiency of different methods, we compare their inference speeds, as shown in Table III. As can be seen from the table, DTM-based methods, namely SDformer and MSDformer, significantly outperform DDPMs-based methods such as Diffusion-TS in terms of inference speed, with a difference of more than 10 times. This is because DDPMs-based methods require multiple denoising steps, leading to longer inference times. Additionally, MSDformer requires more inference time than SDformer. This is because multi-scale modeling, while learning higher-quality discrete representations of time series, also increases the length of the discrete token sequence, thereby increasing the time required for autoregressive generation. Despite the moderate speed cost, MSDformer’s significant quality gains in both short-term and long-term generation (as shown in Section VI-B and Section VI-C) justify this trade-off for most real-world applications. The choice between SDformer and MSDformer can be adapted to specific latency-vs-quality requirements.
VI-E Ablation Studies
@@
Effect of Increasing Codebook Size vs. Number of Scales. Section V-B demonstrates that increasing the number of scales allows for lower distortion compared to increasing the codebook size. Furthermore, we provide an experimental performance comparison of these approaches under different length settings on the fMRI dataset, as shown in Table IV. From the comparison, it is evident that using SDformer with a codebook size of $V=512$ as the baseline, increasing the codebook size to $V=640$ results in performance improvements. However, the performance enhancement is more pronounced and stable when increasing the number of scales (e.g., $V^{(1:K)}=[128,512]$, i.e., MSDformer). These findings align with the theoretical analysis in Section V-B, collectively validating the advantages of multi-scale modeling.
@@
Effect of Similarity-Driven Vector Quantization. Vector quantization critically impacts discrete token learning in time series modeling. Following the finding that similarity-driven strategy outperforms distance-based strategy in time series generation tasks [39], we investigate whether MSDformer retains this advantage by replacing its similarity-driven vector quantization with distance-based variant (”w/o similarity” in Table VI). Results show that MSDformer with similarity-driven vector quantization consistently surpasses distance-based counterparts across all metrics, aligning with SDformer’s finding. this superiority stems from the enhanced capacity of similarity metrics to capture directional features (e.g., trends, periodicity), yielding higher-quality discrete representations that better preserve temporal patterns during modeling.
Effect of Exponential Moving Average for Codebook Update. MSDformer employs exponential moving average (EMA) instead of standard gradient-based codebook updates in VQ-VAE [42]. To validate this design, we compare EMA with conventional methods (”w/o EMA” in Table VI denotes gradient-updated variant). Results reveal consistent performance degradation across metrics when using gradient updates, confirming EMA’s critical role. The EMA strategy ensures stable codebook vector updates while preserving long-term temporal patterns through momentum-based optimization, thereby enhancing model stability and representation fidelity.
Effect of Codebook Reset. MSDformer’s codebook reset mechanism reactivates underused vectors to prevent codebook collapse. Ablation experiments show disabling this technique (”w/o Reset” in Table VI) causes severe performance degradation: Discriminative Score rises from 0.005 to 0.423 and Context-FID increases from 0.009 to 14.664. Codebook utilization plummets to 1.1% without reset, indicating most vectors remain inactive during training. This collapse directly limits representational capacity, confirming the reset’s critical role in maintaining codebook diversity through dynamic vector reactivation.
Effect of Token Type Encoding. MSDformer addresses multi-scale feature confusion through Token Type Encoding (TTE). Ablation experiments reveal disabling TTE (”w/o TTE” in Table VI) causes 40% Discriminative Score degradation and 22.2% Context-FID increase, confirming feature entanglement occurs without explicit scale differentiation. This demonstrates TTE’s critical role in maintaining scale-specific feature spaces during multi-scale temporal modeling.
TABLE VI: Results of ablation study.
@@ Methods
VI-F Visual Analysis
Figure 5: Example of Multi-scale Visualization of Time Series for $K=2$. The first two columns illustrate the multi-scale reconstructions derived from hierarchical representations, specifically showing the coarse-scale reconstruction ($k=1$) and the fine-scale reconstruction ($k=2$). The subsequent two columns display the final reconstruction results and the original data for comparison.
@@ VI-F Visual Analysis
VII Conclusions
This paper introduces a novel DTM-based method, MSDformer, which extends SDformer by incorporating multi-scale modeling, thereby significantly enhancing its ability to handle multi-scale patterns. Additionally, the superiority of the DTM technique in time series generation tasks is demonstrated through the lens of rate-distortion theory. We also provide a theoretical analysis of the advantages of multi-scale modeling. Furthermore, experimental results reveal that DTM-based methods, specifically MSDformer and SDformer, consistently outperform continuous models such as GANs, VAEs, and DDPMs in time series generation (TSG) task. The experiments also highlight the advantages of multi-scale modeling, with MSDformer showing notable improvements over SDformer.
Looking ahead, future work could focus on enhancing scalability by exploring adaptive multi-scale structures that dynamically adjust the number of scales and downsampling rates at different temporal resolutions. The framework could also be extended to more complex scenarios, such as spatiotemporal data generation and multivariate conditional generation, to explore its cross-modal potential. This study not only advances time series generation but also provides valuable insights for other temporal analysis tasks, such as forecasting and anomaly detection.
@@ VII Conclusions
  • In Section I, we give the detailed description of used datasets and the metrics.
  • In Section II, we provide the experiment settings.
  • In Section III, we briefly introduce the baselines.
@@
  • In Section IV, we show more experimental results to verify the effect of increasing codebook size vs. number of scales.
  • In Section VI, we provide the details operations of the architecture in stage 1 and 2.
  • In Section V, we provide the algorithms of training in SDformer and MSDformer.
I Dataset and Metric Details
@@
Dataset. The Stocks dataset consists of Google’s stock price data between 2004 and 2019, with each observation representing a day and containing 6 features. The ETTh dataset includes data obtained from electrical transformers, encompassing load and oil temperature measurements taken every 15 minutes from July 2016 to July 2018. The Energy dataset, a UCI appliance energy prediction dataset, comprises 28 features. The fMRI dataset serves as a benchmark for causal discovery and features realistic simulations of blood-oxygen-level-dependent (BOLD) time series; we chose a simulation with 50 features from the original dataset referring to [26]. The Sines dataset contains 5 features, each generated independently with varying frequencies and phases. The MuJoCo dataset is a multivariate physics simulation time series dataset with 14 features.
@@
Quantitative Metrics. The Discriminative Score measures the distributional similarity between original and synthesized time series data. A binary classifier (e.g., RNN) is trained to distinguish original data (labeled as ”real”) from synthesized data (labeled as ”synthetic”), followed by evaluating its classification accuracy on a test set. The score is derived as the absolute difference between the classifier’s accuracy and the random-guessing baseline (0.5), where a value closer to 0 indicates higher similarity between the two datasets. The Predictive Score evaluates the temporal fidelity of synthesized data by training a forecasting model (e.g., RNN) on synthetic data to predict future time steps and testing its generalization on the original dataset. The score is quantified via the model’s mean absolute error (MAE) on real test samples, where lower values indicate better preservation of the original data’s temporal dynamics. The Context-FID Score evaluates the distributional fidelity of synthesized time series by comparing contextual feature embeddings between original and synthetic datasets. A self-supervised embedding network (e.g., trained via temporal contrastive learning) first extracts context-aware representations from both datasets. The Fréchet distance is then computed between the multivariate Gaussian distributions of these embeddings, characterized by their means and covariance matrices. Lower scores indicate closer alignment in contextual patterns between real and synthetic data.
@@
II Experimental Settings

### v2

@@ 5.1 Theoretical Analysis of Discrete Token Modeling
As an information-theoretic lens, consider a fixed-length time-series window (block) random vector $X_{1:\tau}\sim P_{data}$ and a bounded distortion measure $d:\mathcal{X}^{\tau}\times\tilde{\mathcal{X}}^{\tau}\to\mathbb{R}^{+}$. The minimum achievable rate at distortion level $D$ is characterized by the (block) rate–distortion function:
$R{(D)}=\frac{1}{\tau}\inf_{\begin{subarray}{c}p(\tilde{X}_{1:\tau}\,|\,X_{1:\tau}):\\ \mathbb{E}[d(X_{1:\tau},\tilde{X}_{1:\tau})]\leq D\end{subarray}}I\!\left(X_{1:\tau};\tilde{X}_{1:\tau}\right),$ (23)
where $I(X_{1:\tau};\tilde{X}_{1:\tau})$ denotes the mutual information between the window $X_{1:\tau}$ and its reconstruction $\tilde{X}_{1:\tau}$. $\mathbb{E}[\cdot]$ denotes expectation.
For the time series generation task, let $X_{1:\tau}=(X_{1},\ldots,X_{\tau})\in\mathbb{R}^{\tau\times d}$ denote a random window sampled from the data distribution $P_{data}$ and $\tilde{X}_{1:\tau}$ for its reconstruction through $P_{\mathcal{E},\mathcal{D}}(X_{1:\tau}|\theta_{\mathcal{E}},\theta_{\mathcal{D}})$ of stage 1. The distortion is defined as the window MSE:
@@
In the vector quantization method, the time series data $X_{1:\tau}$ is compressed into a discrete token sequence $y_{1:L}\in\{0,1,\ldots,V-1\}^{L}$ via an encoder $\mathcal{E}$ of stage 1. The rate-distortion control for this method is described in Theorem 2.
@@
$\mathbb{E}_{X_{1:\tau}\sim P_{data},\tilde{X}_{1:\tau}\sim P_{\mathcal{E},\mathcal{D}}(X_{1:\tau}|\theta_{\mathcal{E}},\theta_{\mathcal{D}})}[d(X_{1:\tau},\tilde{X}_{1:\tau})]\leq D.$ (27)
@@
5.2 Theoretical Analysis of Multi-scale Modeling
In Section 5.1, we conclude that increasing the codebook size directly raises the rate, which in turn allows for lower distortion. In MSDformer, the multi-scale modeling, which involves learning multi-scale discrete token representations of time series, indirectly increases the codebook size. To compare the merits of indirectly increasing the codebook size by adding scales versus directly increasing the codebook size on a single scale, this part will conduct a comparative analysis from a theoretical perspective.
@@ 5.2 Theoretical Analysis of Multi-scale Modeling
$\begin{aligned} R_{s}&=\log_{2}(V+V^{\prime})^{L}\\ &=L\log_{2}V\left(1+\frac{V^{\prime}}{V}\right)\\ &=L\log_{2}V+L\log_{2}\left(1+\frac{V^{\prime}}{V}\right)\end{aligned}.$ (31)
@@
$\begin{aligned} R_{m}&=\log_{2}\left(V^{L}{V^{\prime}}^{L^{\prime}}\right)\\ &=L\log_{2}V+L^{\prime}\log_{2}V^{\prime}\end{aligned}.$ (32)
In our experimental setup, the ratios $\frac{V^{\prime}}{V}$ and $\frac{L^{\prime}}{L}$ are constrained to the interval $\left[\frac{1}{4},4\right]$, and they share the same directional relationship (both are either $\geq 1$ or $\leq 1$). Furthermore, $V$ and $V^{\prime}$ are set within the range $2^{7},2^{10}]$. Under these conditions, it is observed that $L\log_{2}\left(1+\frac{V^{\prime}}{V}\right)<L^{\prime}\log_{2}V^{\prime}$, which leads to the conclusion that $R_{m}>R_{s}$ always holds true. Therefore, in general, employing multi-scale modeling is more effective in enhancing the rate compared to directly increasing the codebook size on a single scale, thereby allowing for lower distortion. This theoretically validates the rationality of the MSDformer design.
6 Experiments
@@
6.1 Datasets and Metrics
Datasets. We evaluate the proposed method on six benchmark datasets, including four real-world datasets (Stocks, ETT, Energy, and fMRI) and two simulated datasets (Sines and MuJoCo), which are popular introduced in the previous works [59, 67, 62], and the details of datasets are shown in Appendix A.
Metrics. To quantitatively evaluate the synthetic data, we employ three metrics: 1) The Discriminative Score assesses distributional similarity by measuring how well a classifier can distinguish between real and synthetic samples; 2) The Predictive Score evaluates the practical utility of synthetic data by evaluating the performance of models trained on synthetic samples when tested against real data; 3) The Context-FID Score quantifies sample fidelity through representation learning, specifically comparing the distribution of local temporal features between real and synthetic time series. The details of calculations are shown in Appendix A and all experiments report mean and standard deviation over 5 random seeds.
@@ 6.1 Datasets and Metrics
Implementation Details. For Stage1: we set the codebook sizes and dimensions in $\{128,256,512\}$ and $K$ with $\{2,4,8\}$ across all datasets. For Stage2: we use the decoder-only Transformer with 2 layers, and set $\epsilon$ within $\{0.1,0.3\}$. All experiments are conducted on a single Nvidia V-100 GPU with the AdamW[32] optimizer. More details of our model and configurations are put in Appendix B, duo to page limits.
6.2 Short-term Time Series Generation
TABLE I: Results for all methods on all datasets are averaged over five runs (mean ±.standard deviation).
@@ 6.2 Short-term Time Series Generation
TABLE II: Results of long-term time series generation, the performances are achieved based on 5 runs.
@@
Table I summarizes the performance of each method across all datasets. The results in bold indicate the best performance, while the results with an underline indicate the second-best performance. These results allow us to draw several key observations.
@@
Fig. 5 and Fig. 6 of the Appendix show the comparative visualization results of MSDformer, SDformer, and Diffusion-TS across all datasets. These visualizations depict the distribution of synthesized and real time series data in a two-dimensional space and their probability density distributions. The results indicate that the data generated by MSDformer and SDformer closely resemble real data, highlighting the superior performance of DTM-based methods in time series generation tasks. Furthermore, through a detailed comparison, it can still be observed that the performance of MSDformer is better than that of SDformer. For example, in the Stock dataset, when the x-axis value is around 0.7, the kernel density estimation of the time series data generated by SDformer has a slight deviation from the kernel density estimation of the real data, while MSDformer is significantly closer to the real data at this position. This further illustrates the advantages of multi-scale modeling.
6.3 Long-term Time Series Generation
@@
6.4 Efficiency Analysis
@@
[…]
To validate the practical efficiency of different methods, we compare their inference speeds, as shown in Table III. As can be seen from the table, DTM-based methods, namely SDformer and MSDformer, significantly outperform DDPMs-based methods such as Diffusion-TS in terms of inference speed, with a difference of more than 10 times. This is because DDPMs-based methods require multiple denoising steps, leading to longer inference times. Additionally, MSDformer requires more inference time than SDformer. This is because multi-scale modeling, while learning higher-quality discrete representations of time series, also increases the length of the discrete token sequence, thereby increasing the time required for autoregressive generation. Despite the moderate speed cost, MSDformer’s significant quality gains in both short-term and long-term generation (as shown in Section 6.2 and Section 6.3) justify this trade-off for most real-world applications. The choice between SDformer and MSDformer can be adapted to specific latency-vs-quality requirements.
6.5 Ablation Studies
@@
Effect of Increasing Codebook Size vs. Number of Scales. Section 5.2 demonstrates that increasing the number of scales allows for lower distortion compared to increasing the codebook size. Furthermore, we provide an experimental performance comparison of these approaches under different length settings on the fMRI dataset, as shown in Table IV. From the comparison, it is evident that using SDformer with a codebook size of $V=512$ as the baseline, increasing the codebook size to $V=640$ results in performance improvements. However, the performance enhancement is more pronounced and stable when increasing the number of scales (e.g., $V^{(1:K)}=[128,512]$, i.e., MSDformer). These findings align with the theoretical analysis in Section 5.2, collectively validating the advantages of multi-scale modeling.
@@
Effect of Similarity-Driven Vector Quantization. Vector quantization critically impacts discrete token learning in time series modeling. Following the finding that similarity-driven strategy outperforms distance-based strategy in time series generation tasks [67], we investigate whether MSDformer retains this advantage by replacing its similarity-driven vector quantization with distance-based variant (“w/o similarit” in Table VI). Results show that MSDformer with similarity-driven vector quantization consistently surpasses distance-based counterparts across all metrics, aligning with SDformer’s finding. this superiority stems from the enhanced capacity of similarity metrics to capture directional features (e.g., trends, periodicity), yielding higher-quality discrete representations that better preserve temporal patterns during modeling.
Effect of Exponential Moving Average for Codebook Update. MSDformer employs exponential moving average (EMA) instead of standard gradient-based codebook updates in VQ-VAE [48]. To validate this design, we compare EMA with conventional methods (“w/o EM” in Table VI denotes gradient-updated variant). Results reveal consistent performance degradation across metrics when using gradient updates, confirming EMA’s critical role. The EMA strategy ensures stable codebook vector updates while preserving long-term temporal patterns through momentum-based optimization, thereby enhancing model stability and representation fidelity.
Effect of Codebook Reset. MSDformer’s codebook reset mechanism reactivates underused vectors to prevent codebook collapse. Ablation experiments show disabling this technique (“w/o Reset” in Table VI) causes severe performance degradation: Discriminative Score rises from 0.005 to 0.423 and Context-FID increases from 0.009 to 14.664. Codebook utilization plummets to 1.1% without reset, indicating most vectors remain inactive during training. This collapse directly limits representational capacity, confirming the reset’s critical role in maintaining codebook diversity through dynamic vector reactivation.
Effect of Token Type Encoding. MSDformer addresses multi-scale feature confusion through Token Type Encoding (TTE). Ablation experiments reveal disabling TTE (“w/o TTE” in Table VI) causes 40% Discriminative Score degradation and 22.2% Context-FID increase, confirming feature entanglement occurs without explicit scale differentiation. This demonstrates TTE’s critical role in maintaining scale-specific feature spaces during multi-scale temporal modeling. By explicitly indicating the scale ID of Eqn 17, TTE provides a lightweight yet essential inductive bias for multi-scale discrete time-series modeling, which enables our MSDformer to achieve better generation performance.
TABLE VI: Results of ablation study with the different modeling strategies of vector quantization in stage 1.
@@ Methods
6.6 Visual Analysis
Figure 3: Multi-scale time-series visualization for ($K=2$). The first two columns show hierarchical reconstructions at the coarse (($K=1$)) and fine ($K=2$) scales, respectively. The last two columns present the final reconstruction and the original series for comparison.
@@ 6.6 Visual Analysis
7 Conclusion
This paper introduces a novel DTM-based method, MSDformer, which extends SDformer by incorporating multi-scale modeling, thereby significantly enhancing its ability to handle multi-scale patterns. Additionally, the superiority of the DTM technique in time series generation tasks is demonstrated through the lens of rate-distortion theory. We also provide a theoretical analysis of the advantages of multi-scale modeling. Furthermore, experimental results reveal that DTM-based methods, specifically MSDformer and SDformer, consistently outperform continuous models such as GANs, VAEs, and DDPMs in time series generation (TSG) task. The experiments also highlight the advantages of multi-scale modeling, with MSDformer showing notable improvements over SDformer. Notably, MSDformer exhibits particularly clear gains on long-term generation task, where modeling long-range dependencies and cross-scale structures becomes more challenging. These results indicate that the proposed multi-scale discrete token hierarchy better preserves global dynamics while refining local variations over extended generation length.
Looking ahead, future work could focus on enhancing scalability by exploring adaptive multi-scale structures that dynamically adjust the number of scales and downsampling rates across temporal resolutions. We also plan to explore a unified discrete tokenization strategy for cross-domain generation, enabling a single model to learn transferable token representations across diverse time-series domains and to adapt efficiently to unseen domains. In addition, the framework could be extended to more complex scenarios, such as spatiotemporal data generation and multivariate conditional generation, to further investigate its cross-modal potential. This study not only advances time-series generation but also provides insights for other temporal analysis tasks, such as forecasting and anomaly detection.
@@ 7 Conclusion
  • In Appendix A, we give the detailed description of used datasets and the metrics.
  • In Appendix B, we provide the experiment settings.
  • In Appendix C, we briefly introduce the baselines.
@@
  • In Appendix D, we show more experimental results to verify the effect of increasing codebook size vs. number of scales.
  • In Appendix F, we provide the details operations of the architecture in stage 1 and 2.
  • In Appendix E, we provide the algorithms of training in SDformer and MSDformer.
Appendix A Dataset and Metric Details
@@
Dataset. We conduct extensive evaluations of the proposed MSD framework on six widely-used benchmark datasets for multivariate time-series generation task. For real-world datasets, 1) Stocks [^1] https://finance.yahoo.com/quote/GOOG/history/?p=GOOG dataset consists of Google’s stock price data between 2004 and 2019, with each observation representing a day and containing 6 features, including as features the volume and high, low, opening, closing, and adjusted closing price. 2) ETTh [^2] https://github.com/zhouhaoyi/ETDataset [68] dataset includes data obtained from electrical transformers, encompassing load and oil temperature measurements taken every 15 minutes from July 2016 to July 2018. 3) Energy [^3] https://archive.ics.uci.edu/dataset/374/appliances+energy+prediction dataset is a multivariate appliance energy prediction dataset [5], featuring 28 channels with correlated features. It exhibits noisy periodicity and contains average 10-minute continuous-valued measurements typical of real-world energy consumption data. 4) fMRI [^4] https://www.fmrib.ox.ac.uk/datasets/netsim/ dataset serves as a benchmark for causal discovery and features realistic simulations of blood-oxygen-level-dependent (BOLD) time series; we chose a simulation with 50 features from the original dataset referring to [62].
As for two simulated datasets, 5) Sines [^5] https://github.com/jsyoon0823/TimeGAN is to simulate multivariate sinusoidal sequences of different frequencies $\eta$ and phases $\theta$ providing continuous-valued, periodic, multivariate data where each feature is independent of others. For each dimension $i\in\{1,...,5\},x_{i}(t)=\text{sin}(2\pi\eta t+\theta)$, where $\eta\sim\mathcal{U}[0,1]$ and $\theta\sim\mathcal{U}[-\pi,\pi]$. 6) MuJoCo [^6] https://github.com/google-deepmind/dm_control [47, 41] dataset is a multivariate physics simulation time series dataset, which collects a total of 10,000 simulations of the “Hopper” model from the DeepMind Control Suite and MuJoCo simulator. The position of the body in 2D space is uniformly sampled from the interval $[0,0.5]$. The relative position of the limbs is sampled from the range [2,2][-2,2], and initial velocities are sampled from the interval [5,5][-5,5]. In all, there are 10000 sequences of 100 regularly sampled time points with a feature dimension of 14.
Quantitative Metrics. To be specific,
1). Discriminative Score measures the distributional similarity between original and synthesized time series data. A binary classifier (e.g., RNNs-based) is trained to distinguish original data (labeled as “real”) from synthesized data (labeled as “synthetic”), followed by evaluating its classification accuracy on a test set. The score is derived as the absolute difference between the classifier’s accuracy and the random-guessing baseline (0.5), where a value closer to 0 indicates higher similarity between the two datasets. For a fair comparison, we follow the standard TimeGAN and Diffusion-ts [59, 62] style post-hoc evaluation protocol: we train an off-the-shelf RNN-based binary classifier to discriminate real sequences from generated sequences on a held-out test set, and compute the discriminative score as $|\text{Acc}-0.5|$ (lower is better). And the classifier is a GRU-based sequence classifier consisting of a GRU structure followed by a fully-connected sigmoid head. Specifically, given the input sequences $x\in\mathbb{R}^{T\times d}$ (Ground truths and generation results), where $T$ is the generation length and $d$ is the dimension of dataset, we run a GRU model and map it to a scalar logit via a linear layer; the probability is obtained by a sigmoid. Also, within the training setup, we construct a balanced dataset by labeling each original sequence as real (1) and each generated sequence as fake (0). The classifier is trained using binary cross-entropy with the Adam optimizer for 2000 iterations and batch size 128, where each mini-batch contains samples from both real and generated training sets, which follows the experimental settings of Diffusion-ts [^7] https://github.com/Y-debug-sys/Diffusion-TS/tree/main. We evaluate accuracy on the held-out test split and report the discriminative score.
2). Predictive Score evaluates the temporal fidelity of synthesized data by training a post-hoc one-step-ahead sequence predictor (e.g., RNNs-based) on synthetic data to predict future time steps and testing its generalization on the original dataset. The score is quantified via the model’s mean absolute error (MAE) on real test samples, where lower values indicate better preservation of the original data’s temporal dynamics and we also follow the previous works’s experimental settings for fair evaluation. Concretely, given a input sequence $x_{1:T}\in\mathbb{R}^{T\times d}$, the predictor takes the prefix $x_{1:T-1}$ and predicts the next-step sequence $x_{2:T}$. In our implementation, the input context length is (T-1) time steps (all but the last step of each window), and the prediction target is also (T-1) steps, corresponding to a one-step-shifted forecasting objective. Thus, the predictive score does not evaluate long-horizon forecasting; it evaluates one-step-ahead prediction across the entire window following TimeGAN’s post-hoc predictive score protocol. The predictor is trained with L1 loss (MAE) using Adam for 5000 iterations with batch size 128 and the structure is aslo a GRU-based RNN model. Each training iteration samples a mini-batch of synthetic sequences, constructs input–target pairs ($x_{1:T-1},x_{2:T}$), and performs one optimizer step. After training, we evaluate the trained predictor on real sequences and average MAE over all sequences to obtain the predictive score. Overall, it trains a one-step-ahead predictor (shift-by-one) on synthetic data and reports the MAE on real data over all time steps in the generation window.
@@
3). Context-FID Score evaluates the distributional fidelity of synthesized time series by comparing contextual feature embeddings between original and synthetic datasets. A self-supervised embedding network (e.g., trained via temporal contrastive learning) first extracts context-aware representations from both datasets. The Fréchet distance is then computed between the multivariate Gaussian distributions of these embeddings, characterized by their means and covariance matrices. And the Context-FID Score is to replace the Inception model of the original Frechet Inception distance (FID) [23] with a time series representation learning method called TS2Vec [63]. Lower scores indicate closer alignment in contextual patterns between real and synthetic data. Specifically, we first sample synthetic time series and real-time series respectively. Then we compute the FID score of the representation after encoding them with a pre-trained TS2Vec model across each dataset, and each embedding is computed from the entire generation sequence, not from a randomly selected sub-range. Details are shown in Fig. 4.
@@
Input: $x\in\mathbb{R}^{B\times T\times D}$
$\Downarrow$
NaN mask: $m_{\text{nan}}\in\{0,1\}^{B\times T}$, set $x[\neg m_{\text{nan}}]\leftarrow 0$
$\Downarrow$
Linear projection: $\mathrm{Linear}(D\!\rightarrow\!H)$, $h^{(0)}\in\mathbb{R}^{B\times T\times H}$
$\Downarrow$
Time masking (train): sample $m\in\{0,1\}^{B\times T}$, apply $m\leftarrow m\wedge m_{\text{nan}}$, set $h^{(0)}[\neg m]\leftarrow 0$
$\Downarrow$
Transpose: $B\times T\times H\rightarrow B\times H\times T$
$\Downarrow$
Dilated Conv Encoder: residual ConvBlocks with dilation $2^{i}$ and SamePadConv (length preserved), output $\hat{h}^{(L)}\in\mathbb{R}^{B\times C_{\text{out}}\times T}$
$\Downarrow$
Dropout: $p=0.1$
$\Downarrow$
Transpose back: $B\times C_{\text{out}}\times T\rightarrow B\times T\times C_{\text{out}}$
$\Downarrow$
Output: $z\in\mathbb{R}^{B\times T\times C_{\text{out}}}$, where $C_{\text{out}}=\text{output\_dims}$
Figure 4: Details of the TS2Vec encoder.
Appendix B Experimental Settings